OpenStudy (bluefire113):

how do I simplify this question? ((1/xy)-1)/((1/x^2)-(y/x)) the answer is x/y, but i would like to know the steps of solving this problem.

OpenStudy (seratul):

If you are to rewrite it at first, you should get \[\frac{ y/x-1 }{ 1/x ^{2}-y/x }\] This is simply by multiplying by 1 and removing the parenthesis.

OpenStudy (seratul):

I don't really know how I can write this all using the equation tap on here, but you should eventually get \[\frac{ -x^2+xy }{-xy +1 }\]

OpenStudy (bluefire113):

can you explain why y/x got changed into y/x

OpenStudy (bluefire113):

sorry i mean how did 1/xy get changed into y/x

OpenStudy (seratul):

It can be rewritten as y *1 /x which is y/x

OpenStudy (bluefire113):

would this be correct: (y/x-1) * (x^2-x/y)

jhonyy9 (jhonyy9):

1 --- -1 = (1-xy)/xy xy 1 y 1 - xy ---- - ---- = ---------- x^2 x x^2 1 -xy ------- xy 1 -xy x^2 x ----------- = -------- * --------- = ---- 1 -xy xy 1 -xy y ---------- x^2 hope helped

OpenStudy (bluefire113):

Thanks!

jhonyy9 (jhonyy9):

np welcome was my pleasure do you understand the way how you get the right answer ?

OpenStudy (bluefire113):

Yes thanks!

jhonyy9 (jhonyy9):

ok good luck and hope i can help you anytime just tag me than i m here online bye bye